Hi there Savannah State University 3. River rafters often fill their rafts in the chill of the morning. rafting down the cool While river their rafts stay cool. However, when they pull out for lunch and take their rafts part- way out of the river, their rafts can warm up quickly. If a raft warms up too much, they can expand to the point of explosion. Needless to say, a raft with an exploded tube is not very useful in the middle of a wilderness river trip. a. What can rafters do to prevent raft explosions? b, Assume that the raft was filled in the morning when it was 50 F. The initial volume of one of the rafts tubes was 19 ft At lunch on a Grand Canyon expedition, the rafts tube popped when it reached 110 F. What was the volume of air, in ft, when the tube popped?
Answer #1

According to Charles law for a given amount of gas at constant pressure Volume \propto Temperature.

If the volume of air in the raft is decreased when they take out for lunch raft won't explode as now it will take time for air to expand .

B. V1/T1 = V2/T2 ( Temperatures need to convert in Kelvin)

19/283.15 = V2/316.15

V2 = 21.21 ft3